# Goal-Reaching Trajectory Design Near Danger with Piecewise Affine Reach-avoid Computation

### v3

@@ VII Hardware Demonstration
Thus far, PARC has only been shown to work in simulation. In reality, the tracking model will have model mismatch with the actual hardware due to disturbances and perception errors. As such, we now present preliminary results for PARC’s reach-avoid guarantees applied on mobile robot hardware navigating a narrow gap, as shown in Fig. 13. Specifically, we use a TurtleBot3 Burger Model [grossimplementation] differential drive robot running ROS 2 [macenski2022robot].
@@ VIII Conclusion
Limitations: PARC has several key limitations. First and foremost, it requires an assumption that the data coverage of sampled tracking error is large enough. While this assumption holds in practice (over hundreds of collision-free trials planned in this paper), it means that further mathematical work is necessary for true safety guarantees. Second, if the planning of the robot takes place at the joint level, which is common for manipulator and legged robots, augmentation of the obstacles with tracking error is not straight-forward. Though this has been successfully done in other set-based planner-tracker reach-avoid methods such as ARMTD [holmes2020reachable] and ARMOUR [michaux2023can] by expressing H-polytopes in configuration space as polynomial zonotopes [kochdumper2020reachability] in workspace, we opted to leave its extension to PARC as future work, as the focus of this paper is just computing the BRAS in the first place. Third, although PARC is faster than compared methods, it is still too slow for real-time; the present implementation has not been parallelized and may benefit from alternative representations to H-polytopes. Fourth, PARC has only been compared on hand-crafted examples of goal reaching near danger, and hardware validation has only been done on a simple robot. The experiments have only considered tracking error, with the hardware validation considering error from localization drift of the IMU and encoders in addition. It is critical to validate the method across randomly-generated, realistic environments on a variety of hardware with multiple significant sources of uncertainties, such as perception error. To this end, the authors are preparing a small robotic car to perform drift parking experiments in the coming months.
@@
We now discuss methods that guarantee safety (avoiding dangerous states) and liveness (reaching a goal) for control systems and robot motion planning. In particular, we discuss optimal control methods, safety filters, and planner-tracker frameworks. We also note connections between our approach and sampling-based motion planning. Finally, we discuss piecewise affine (PWA) system models, which we leverage to overcome limitations of conservativeness and computational tractability for providing safety and liveness guarantees with state-of-the-art methods.
@@
Many safety methods focus on synthesizing or analyzing controllers. There has been extensive development of techniques such as Control Barrier Functions (CBFs) [ames2016control, agrawal2017discrete, rosolia2022unified, castaneda2022probabilistic], Hamilton-Jacobi (HJ) reachability [herbert2021scalable, seo2022real], sums-of-squares reachability [majumdar2017funnel, tedrake2010lqr], robust model predictive control (MPC) [rakovic2012parameterized, villanueva2017robust, wabersich2021predictive], and graphs of convex sets (GCS) [marcucci2023motion, cohn2023non]. Yet, only a minority of these methods address the dual objectives of liveness and safety [hsu2021safety, so2023solving, borquez2023safety, fridovich2019safely]. Some methods seek a balance between performance and safety by implicitly integrating each constraint into the cost function of optimal control problem, solved through methods like MPC [williams2016aggressive] or reinforcement learning [srinivasan2020learning], but can fail to offer hard satisfaction of such constraints.
@@
Two representative approaches are FaSTrack [chen2021fastrack] and Reachability-Based Trajectory Design (RTD) [kousik2019safe, kousik2020bridging]. FaSTrack employs Hamilton-Jacobi (HJ) reachability to precompute a worst-case error bound between the planning and tracking models, then dilates obstacles by this bound to ensure safe planning online. RTD precomputes a forward reachable set (FRS) of a parameterized set of planning model trajectories and associated (time-varying) tracking error, then uses the FRS to find safe plans at runtime. We compare our proposed method against both FaSTrack and RTD.
@@
In this paper, instead of sampling to find paths or trajectories, we directly approximate a continuum of all feasible trajectories through a narrow gap. That said, we do not view our approach as competing with sampling-based motion planning. By contrast, our PARC planner is complementary to sampling in three ways. First, PARC requires expert plans that seed our reachability analysis; in this paper, we generate the plans by approximating the solution to a two-point boundary value problem, but we could leverage a sampling approach instead. Second, as we show in our experiments, PARC is only necessary when a goal is specifically near obstacles; a sampling-based approach that ignores dynamics could be used to find goals and identify these scenarios (e.g., if the sampling-based approach fails, then one should switch to using PARC). Third, our approach can be used to robustify a sampling-based approach by computing safety funnels around a sampled plan, similar to LQR trees [tedrake2010lqr] or funnel libraries [majumdar2017funnel].
@@
In this paper, we propose a parameterized, time-variant piecewise-affine (PWA) system as our planning model. PWA systems, which evolve through different affine models based on state regions (modes), adeptly capture nonlinear dynamics. Their capacity for modeling non-smooth dynamics through hybridization is well-documented [asarin2007hybridization, dang2010accurate] and they are prevalently used to approximate complex nonlinear hybrid systems, including legged locomotion [marcucci2017approximate], traffic systems [mehr2017stochastic], and neural network [vincent2021reachable].
@@
which can be computed as one H-polytope by solving a linear program (LP) [herceg2013multi]. The Minkowski sum essentially adds a “buffer” to the object being summed, which we use to pad the obstacles to account for tracking error in our reachability analysis.
@@
$\begin{split}Q=\{\mathbf{x}_{\mathrm{\textnormal{ETI}}}&\in\overline{\Omega}_{\mathrm{\textnormal{ETI}}}\mid~{}\exists(\mathbf{x}_{\mathrm{\textnormal{nonETI}}},\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}},\gamma)\in\overline{\Omega}_{\mathrm{\textnormal{nonETI}}}\times\mathcal{O}_{\mathrm{\textnormal{ETI}}}\times[0,1]\quad\\ &\mathrm{\textnormal{s.t.}}\quad\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\gamma(\mathbf{C}_{1}\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\mathbf{x}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\}\end{split}$ (70)
@@
$\mathbf{y}+\mathbf{C}_{1}\mathbf{y}+$ $\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}$ (77)
$\begin{split}=\hat{\mathbf{x}}&{}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})\\ &+\mathbf{C}_{1}(\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}))\\ &+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}\\ \end{split}$ (78)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}^{2}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}-\mathbf{C}_{1}\mathbf{d}_{1}$ (79)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})$ (80)
$=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$ (81)
The equation (81) follows from (80) using ETI Assumption 2 and the result of Lemma 12, namely:
$\begin{split}\mathbf{C}_{X_{\mathrm{\textnormal{ETI}}}}{(\mathbf{C}_{s_{t},t}\mathbf{x}+\mathbf{d}_{s_{t},t})}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}=\mathbf{C}_{1}(\mathbf{C}_{1}\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\mathbf{x}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})&=0\\ \forall\mathbf{x}&\in X\end{split}$ (82)
@@
-E Experiment Details
-E1 Near-Hover Quadrotor in 3-D
@@
where the states $(p_{x},p_{y},p_{z})$ denote the position, $(v_{x},v_{y},v_{z})$ are velocities, $(\theta_{x},\theta_{y})$ are pitch and roll, and $(\omega_{x},\omega_{y})$ are pitch and roll rates. The control input $\mathbf{u}=(\alpha_{x},\alpha_{y},\alpha_{z})$ is the desired pitch and roll $(\alpha_{x},\alpha_{y})$ and the vertical thrust $\alpha_{z}$. The model parameters $(d_{0},d_{1},n_{0},k_{T})$ and control input bound $U$ are referenced from FaSTrack [chen2021fastrack]. A simple Linear Quadratic Regulator (LQR) linearized around the tracking error $e=0$ is employed for the feedback controller.
-E2 General Quadrotor in 3-D
@@
Further, we note that for each of the low-dimensional model, the intermediate avoid sets without tracking error $\overline{\Lambda}_{i,t,t,x}$, $\overline{\Lambda}_{i,t,t,y}$, and $\overline{\Lambda}_{i,t,t,z}$ and hence the avoid set $\Lambda$ can be computed without using convex hull or projection according to the following theorem:
Theorem 23 (Intermediate Avoid Set w/o Convex Hull or Projection).
$k_{i,\mathrm{\textnormal{min}}}$ $=\min_{\mathbf{k}}\{(\mathbf{k})_{i:i}\ |\ \mathbf{k}\in K\},$ (91)
$k_{i,\mathrm{\textnormal{max}}}$ $=\max_{\mathbf{k}}\{(\mathbf{k})_{i:i}\ |\ \mathbf{k}\in K\},$ (92)
for $i=1,\cdots,{n}_{k}$, then
can replace (29) to fulfill the condition in (30).
Proof.
@@
FaSTrack is a planner-tracker framework that precomputes a uniform Tracking Error Bound (TEB) (maximum possible tracking error) using Hamilton-Jacobi (HJ) reachability analysis. Obstacles are padded by the TEB; then, as long as the planner only plans outside the padded obstacles the robot is safe. During the TEB computation, FaSTrack also synthesizes a controller to guarantee the bound. We use RRT as the planning model as per [chen2021fastrack].
@@
-F Additional Experiment Details
-F1 Impact of Planner-Tracker Cooperation
@@
The environment is a $[0,10]\times[-10,10]\times[0,10]$ cuboid (all lengths are in meters). The goal $\overline{P_{\mathrm{\textnormal{G}}}}$ is a cube centered at $(p_{x},p_{y},p_{z})=(9,0,5)$ with a side of $1$unitm. Obstacles $\overline{\mathcal{O}}$ include the floor, ceiling, and walls, and two additional cuboid obstacles of $[6.5,7.5]\times[-8.2,-\Delta w/2]\times[0,10]$ and $[6.5,7.5]\times[\Delta w/2,8.2]\times[0,10]$ with varying $\Delta w$ that denotes the gap between two obstacles. These are shown in Fig. 7 with a gap width $\Delta w=1.8$. The obstacles are padded to accommodate the quadrotor’s body, approximated by a cuboid of $0.54\times 0.54\times 0.05$ using the specifications of Hummingbird [dong2015development]. The planning horizon is $t_{\mathrm{\textnormal{f}}}=10$units with time discretization $\Delta t=0.1$units each. The frequency of the tracking controller is $1000$unitHz for all methods.
-F2 Impact of Tighter Planning Reachability
@@
The general 3-D quadrotor environment is similar to that of the 10-D near-hover quadrotor environment but with a narrower gap to the goal. The obstacles are defined as $\overline{\mathcal{O}}_{1}=[3.23,6.77]\times[0.23,9.27]\times[0.73,9.27]$, $\overline{\mathcal{O}}_{2}=[3.23,6.77]\times[-9.27,-0.23]\times[0.73,9.27]$ (after accounting for an overapproximated volume of the quadrotor), that is, two walls with a clearance of 0.46, the goal region is defined as $\overline{P_{\mathrm{\textnormal{G}}}}=[7.44,9.56]\times[-1.06,1.06]\times[3.94,6.06]$, and the domains are defined as $P=[0,10]\times[-10,10]\times[0,10]$ and $K=[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]$. The quadrotor starts stationary at $p_{x}=1,p_{y}=0,p_{z}=5$. This means the quadrotor begins at the side of the wall opposite to the goal region, but the $p_{y}$ coordinate is at the middle of the gap.
-F3 Impact of Proper Use of Time-varying Tracking Error
@@
In the 3-D quadrotor environment, similar to the one described in Section V-C, we evaluate two scenarios differentiated by obstacle gap widths. The obstacles are specified for both scenarios as $\overline{\mathcal{O}}_{1}=[5.75,7.25]\times[\triangle w/2,10]\times[1,9]$ and $\overline{\mathcal{O}}_{2}=[5.75,7.25]\times[-10,-\triangle w/2]\times[1,9]$, using $\triangle w=0.85$ for the narrow gap and $\triangle w=3.0$ for the wide gap. The goal region is set as $\overline{P_{\mathrm{\textnormal{G}}}}=[7,9]\times[-1,1]\times[4,6]$, with planning domains defined by $P=[0,10]\times[-10,10]\times[0,10]$. The trajectory parameter space, $K=[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]$, was selected to simplify the tuning of a funnel-based feedback controller. Initial states for simulations are sampled from a $(25,25,3)$ grid spanning $[0.1,4.8]\times[-9.9,9.9]\times[3,7]$. Due to uniform obstacle distribution in the $p_{z}$ direction, sampling is denser in the $p_{x},p_{y}$ plane, since we expect varying initial states in $p_{z}$ direction would not be a significant influence. Each simulation starts with the quadrotor stationary, following reference trajectories determined by each methodology.

### v4

@@ VII Hardware Demonstration
Thus far, PARC has only been shown to work in simulation. In reality, the tracking model will have model mismatch with the actual hardware due to disturbances and perception errors. As such, we now present preliminary results for PARC’s reach-avoid guarantees applied on mobile robot hardware navigating a narrow gap, as shown in Fig. 13. Specifically, we use a TurtleBot3 Burger Model [grossimplementation] differential drive robot running Robot Operating System (ROS) 2 [macenski2022robot].
@@ VIII Conclusion
Limitations: PARC has several key limitations.
First and foremost, it requires an assumption that the data coverage of sampled tracking error is large enough. While this assumption holds in practice (over hundreds of collision-free trials planned in this paper), it means that further mathematical work is necessary for true safety guarantees.
Second, if the planning of the robot takes place at the joint level, which is common for manipulator and legged robots, augmentation of the obstacles with tracking error is not straight-forward. Though this has been successfully done in other set-based planner-tracker reach-avoid methods such as Autonomous Reachability-based Manipulator Trajectory Design (ARMTD) [holmes2020reachable] and Autonomous Robust Manipulation via Optimization with Uncertainty-aware Reachability (ARMOUR) [michaux2023can] by expressing H-polytopes in configuration space as polynomial zonotopes [kochdumper2020reachability] in workspace, we opted to leave its extension to PARC as future work, as the focus of this paper is just computing the BRAS in the first place.
Third, although PARC is faster than compared methods, it is still too slow for real-time; the present implementation has not been parallelized and may benefit from alternative representations to H-polytopes.
Fourth, depending on the choice of planning model, PARC could be difficult to implement on long-horizon problem settings. This is because the assumption of having obtained an expert plan with constant trajectory parameters that connects the initial position to the goal set (Assumption 9), if such plan exists at all, gets increasingly difficult to fulfill the longer the planning horizon is [bonalli2019gusto, lofberg2004yalmip]. Though a good choice of planning model (e.g. time-variant affine systems) would alleviate this problem, we restrict the problem setups in this paper to short-horizon (i.e. near danger) settings to avoid overclaiming the capabilities of PARC.
Lastly, PARC has only been compared on hand-crafted examples of goal-reaching near danger, and hardware validation has only been done on a simple robot. The experiments have only considered tracking error, with the hardware validation considering error from localization drift of the IMU and encoders in addition. It is critical to validate the method across randomly-generated, realistic environments on a variety of hardware with multiple significant sources of uncertainties, such as perception error. To this end, the authors are preparing a small robotic car to perform drift parking experiments.
Acknolwedgments
We are grateful to the anonymous reviewers and editor’s suggestions for additional comparisons and presentation. We would also like to thank Trey Weber for assistance with the drifting dynamics, and Christos Verginis for assistance with implementing KDF.
@@
We now discuss methods that guarantee safety (avoiding dangerous states) and liveness (reaching a goal) for control systems and robot motion planning. In particular, we discuss optimal control methods, safety filters, and planner-tracker frameworks. We also note connections between our approach and sampling-based motion planning. Finally, we discuss PWA system models, which we leverage to overcome limitations of conservativeness and computational tractability for providing safety and liveness guarantees with state-of-the-art methods.
@@
Many safety methods focus on synthesizing or analyzing controllers. There has been extensive development of techniques such as CBFs [ames2016control, agrawal2017discrete, rosolia2022unified, castaneda2022probabilistic], Hamilton-Jacobi (HJ) reachability [herbert2021scalable, seo2022real], sums-of-squares (SOS) reachability [majumdar2017funnel, tedrake2010lqr], robust MPC [rakovic2012parameterized, villanueva2017robust, wabersich2021predictive], and graphs of convex sets (GCS) [marcucci2023motion, cohn2023non]. Yet, only a minority of these methods address the dual objectives of liveness and safety [hsu2021safety, so2023solving, borquez2023safety, fridovich2019safely]. Some methods seek a balance between performance and safety by implicitly integrating each constraint into the cost function of optimal control problem, solved through methods like MPC [williams2016aggressive] or reinforcement learning (RL) [srinivasan2020learning], but can fail to offer hard satisfaction of such constraints.
@@
Two representative approaches are FaSTrack [chen2021fastrack] and RTD [kousik2019safe, kousik2020bridging]. FaSTrack employs HJ reachability to precompute a worst-case error bound between the planning and tracking models, then dilates obstacles by this bound to ensure safe planning online. RTD precomputes a forward reachable set (FRS) of a parameterized set of planning model trajectories and associated (time-varying) tracking error, then uses the FRS to find safe plans at runtime. We compare our proposed method against both FaSTrack and RTD.
@@
In this paper, instead of sampling to find paths or trajectories, we directly approximate a continuum of all feasible trajectories through a narrow gap. That said, we do not view our approach as competing with sampling-based motion planning. By contrast, our PARC planner is complementary to sampling in three ways. First, PARC requires expert plans that seed our reachability analysis; in this paper, we generate the plans by approximating the solution to a two-point boundary value problem, but we could leverage a sampling approach instead. Second, as we show in our experiments, PARC is only necessary when a goal is specifically near obstacles; a sampling-based approach that ignores dynamics could be used to find goals and identify these scenarios (e.g., if the sampling-based approach fails, then one should switch to using PARC). Third, our approach can be used to robustify a sampling-based approach by computing safety funnels around a sampled plan, similar to linear quadratic regulator (LQR) trees [tedrake2010lqr] or funnel libraries [majumdar2017funnel].
@@
In this paper, we propose a parameterized, time-variant PWA system as our planning model. PWA systems, which evolve through different affine models based on state regions (modes), adeptly capture nonlinear dynamics. Their capacity for modeling non-smooth dynamics through hybridization is well-documented [asarin2007hybridization, dang2010accurate] and they are prevalently used to approximate complex nonlinear hybrid systems, including legged locomotion [marcucci2017approximate], traffic systems [mehr2017stochastic], and neural network [vincent2021reachable].
@@
which can be computed as one H-polytope by solving an LP [herceg2013multi]. The Minkowski sum essentially adds a “buffer” to the object being summed, which we use to pad the obstacles to account for tracking error in our reachability analysis.
@@
$\begin{split}Q=&\{\mathbf{x}_{\mathrm{\textnormal{ETI}}}\in\overline{\Omega}_{\mathrm{\textnormal{ETI}}}\mid~{}\exists(\mathbf{x}_{\mathrm{\textnormal{nonETI}}},\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}},\gamma)\in\overline{\Omega}_{\mathrm{\textnormal{nonETI}}}\times\mathcal{O}_{\mathrm{\textnormal{ETI}}}\\ &\times[0,1]\ \mathrm{\textnormal{s.t.}}\ \mathbf{x}_{\mathrm{\textnormal{ETI}}}+\gamma(\mathbf{C}_{1}\mathbf{x}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\mathbf{x}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})=\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\}\end{split}$ (66)
@@
$\mathbf{y}+\mathbf{C}_{1}\mathbf{y}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}$
$=$ $\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})$
$+\mathbf{C}_{1}(\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}))+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1}$ (73a)
$=$ $\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}-\mathbf{C}_{1}(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})$ (73b)
$=$ $\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}\in\mathcal{O}_{\mathrm{\textnormal{ETI}}}$ (73c)
The equation (73c) follows from (73b) using ETI Assumption 2 and the result of Lemma 11 (i.e. (25)) namely:
$\begin{split}&\left\|\mathbf{C}_{1}(\mathbf{C}_{1}\hat{\mathbf{x}}_{\mathrm{\textnormal{ETI}}}+\mathbf{C}_{2}\tilde{\mathbf{x}}_{\mathrm{\textnormal{nonETI}}}+\mathbf{d}_{1})\right\|_{2}\\ =&\left\|\hat{\mathbf{C}}_{s_{t},t}^{X_{\mathrm{\textnormal{ETI}}}}{(\hat{\mathbf{C}}_{s_{t},t}\mathbf{x}+\mathbf{d}_{s_{t},t})}_{1:{{n}_{x}}_{\mathrm{\textnormal{ETI}}}}\right\|_{2}\\ =&0\end{split}$ (74)
@@
-E Alternate Methods for Computing the Avoid Set
Here, we present two methods to compute the avoid set without tracking error as alternatives to Corollary 14 and Theorem 12. To account for tracking error, the goal set and the obstacles can be augmented or subtracted with the corresponding error sets as in Section IV-E2.
-E1 Grid-Based Avoid Set
Recall that, per (8), we model the continuous-time behavior as straight line segments between the discrete states of the PWA system. Intuitively, the avoid set should include all trajectories where the straight lines connecting the PWA planning model’s discrete states collide with any obstacles. Thus, we can establish a necessary condition for collision to occur between two discrete timesteps with the following Proposition:
Proposition 21 (Unsafe Plans w/o Tracking Error).
[…]
then
Proof.
By the definition of convex-hull, we have
[…]
$\mathbf{x}(t),\mathbf{x}(t+\Delta t)\in\overline{\Omega}_{t}\cup\overline{\Omega}_{t+\Delta t}\},$ (77a)
[…]
Thus (75) necessarily implies (76). ∎
One can exploit Proposition 21 to compute the avoid set by gridding up the goal set and checking for each grid if any of their intermediate BRSs satisfies (76) with any of the obstacles. The union of all grids that fulfill such condition can then be taken as an over-approximation of the true avoid set without tracking error. Alternatively, one can also grid up the reach set and check if their FRS [herceg2013multi] intersects with any of the obstacles instead.
Since this method does not require ETI assumptions on the PWA planning model, one can show that its approximation of the true avoid set gets tighter as the grid gets smaller. In practice, we found that computing avoid set in this manner is significantly slower than Theorem 12 for a desirable grid size. Thus, we have only used it in Fig. 4 to show the extend of overapproximation of Theorem 12.
-E2 Avoid Set w/o Convex Hull or Projection
In PWA planning models with specific structures, the intermediate avoid sets without tracking error $\overline{\Lambda}_{i,t,t}$, and hence the avoid set $\Lambda$ can sometimes be computed without using convex hull or projection, significantly speeding up and reducing the conservativeness of the approximation:
Theorem 22 (Intermediate Avoid Set w/o Convex Hull or Projection).
$k_{i,\mathrm{\textnormal{min}}}$ $=\min_{\mathbf{k}}\{\mathbf{k}_{i}\ |\ \mathbf{k}\in K\},$ (80)
$k_{i,\mathrm{\textnormal{max}}}$ $=\max_{\mathbf{k}}\{\mathbf{k}_{i}\ |\ \mathbf{k}\in K\},$ (81)
for $i=1,\cdots,{n}_{k}$, where $\mathbf{k}_{i}$ is the $i^{\mathrm{\textnormal{th}}}$ element of $\mathbf{k}$, then
can replace (27) to fulfill the condition in (28).
Proof.
In this paper, the PWA planning models in Section V-C, Section V-D, and Appendix -H fulfill this condition at all timesteps.
-F Experiment Details
-F1 Near-Hover Quadrotor in 3-D
@@
where the states $(p_{x},p_{y},p_{z})$ denote the position, $(v_{x},v_{y},v_{z})$ are velocities, $(\theta_{x},\theta_{y})$ are pitch and roll, and $(\omega_{x},\omega_{y})$ are pitch and roll rates. The control input $\mathbf{u}=(\alpha_{x},\alpha_{y},\alpha_{z})$ is the desired pitch and roll $(\alpha_{x},\alpha_{y})$ and the vertical thrust $\alpha_{z}$. The model parameters $(d_{0},d_{1},n_{0},k_{T})$ and control input bound $U$ are referenced from FaSTrack [chen2021fastrack]. A simple LQR linearized around the tracking error $e=0$ is employed for the feedback controller.
-F2 General Quadrotor in 3-D
@@
Further, we note that for each of the low-dimensional model, the intermediate avoid sets without tracking error $\overline{\Lambda}_{i,t,t,x}$, $\overline{\Lambda}_{i,t,t,y}$, and $\overline{\Lambda}_{i,t,t,z}$ and hence the avoid set $\Lambda$ can be computed without using convex hull or projection according to Theorem 22.
@@
FaSTrack is a planner-tracker framework that precomputes a uniform TEB (maximum possible tracking error) using HJ reachability analysis. Obstacles are padded by the TEB; then, as long as the planner only plans outside the padded obstacles the robot is safe. During the TEB computation, FaSTrack also synthesizes a controller to guarantee the bound. We use RRT as the planning model as per [chen2021fastrack].
@@
KDF
KDF combines sampling-based motion planning with funnel-based feedback control [verginis2022kdf]. This method keeps the tracking error within funnels defined by a performance function which decreases exponentially over time [bechlioulis2014low]. These funnels help higher-level motion planners by incorporating maximum funnel values to expand the free space for planning. The implementation details are described in Appendix -G3.
-G Additional Experiment Details
-G1 Impact of Planner-Tracker Cooperation
@@
The environment is a $[0,10]\times[-10,10]\times[0,10]$ cuboid (all lengths are in meters). The goal $\overline{P_{\mathrm{\textnormal{G}}}}$ is a cube centered at $(p_{x},p_{y},p_{z})=(9,0,5)$ with a side of $1$unitm. Obstacles $\overline{\mathcal{O}}$ include the floor, ceiling, and walls, and two additional cuboid obstacles of $[6.5,7.5]\times[-8.2,-\Delta w/2]\times[0,10]$ and $[6.5,7.5]\times[\Delta w/2,8.2]\times[0,10]$ with varying $\Delta w$ that denotes the gap between two obstacles. These are shown in Fig. 7 with a gap width $\Delta w=1.8$. The obstacles are padded to accommodate the quadrotor’s body, approximated by a cuboid of $0.54\times 0.54\times 0.05$ using the specifications of Hummingbird [dong2015development]. The planning horizon is $t_{\mathrm{\textnormal{f}}}=10$units with time discretization $\Delta t=0.1$units each. The frequency of the tracking controller is $1,000$unitHz for all methods.
-G2 Impact of Tighter Planning Reachability
@@
The general 3-D quadrotor environment is similar to that of the 10-D near-hover quadrotor environment but with a narrower gap to the goal. The obstacles are defined as $\overline{\mathcal{O}}_{1}=[3.23,6.77]\times[0.23,9.27]\times[0.73,9.27]$, $\overline{\mathcal{O}}_{2}=[3.23,6.77]\times[-9.27,-0.23]\times[0.73,9.27]$ (after accounting for an overapproximated volume of the quadrotor), that is, two walls with a clearance of 0.46, the goal region is defined as $\overline{P_{\mathrm{\textnormal{G}}}}=[7.44,9.56]\times[-1.06,1.06]\times[3.94,6.06]$, and the domains are defined as $P=[0,10]\times[-10,10]\times[0,10]$ and $K=[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]\times[-5.25,5.25]\times[-10,10]\times[-5.25,5.25]$. Initial positions for simulations are sampled from a $(15,15,3)$ grid spanning $[0.1,4.8]\times[-9.9,9.9]\times[3,7]$, all with zero initial velocity.
-G3 Impact of Proper Use of Time-varying Tracking Error
@@
In the 3-D quadrotor environment, similar to the one described in Section V-C, we evaluate two scenarios differentiated by obstacle gap widths. The obstacles are specified for both scenarios as $\overline{\mathcal{O}}_{1}=[5.75,7.25]\times[\triangle w/2,10]\times[1,9]$ and $\overline{\mathcal{O}}_{2}=[5.75,7.25]\times[-10,-\triangle w/2]\times[1,9]$, using $\triangle w=0.85$ for the narrow gap and $\triangle w=3.0$ for the wide gap. The goal region is set as $\overline{P_{\mathrm{\textnormal{G}}}}=[7,9]\times[-1,1]\times[4,6]$, with planning domains defined by $P=[0,10]\times[-10,10]\times[0,10]$. The trajectory parameter space, $K=[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]\times[-1.25,1.25]\times[-1.5,1.5]\times[-2,2]$, was selected to simplify the tuning of a funnel-based feedback controller. Initial positions for simulations are sampled from a $(25,25,3)$ grid spanning $[0.1,4.8]\times[-9.9,9.9]\times[3,7]$. Due to uniform obstacle distribution in the $p_{z}$ direction, sampling is denser in the $p_{x},p_{y}$ plane, since we expect varying initial states in $p_{z}$ direction would not be a significant influence. Each simulation starts with the quadrotor stationary, following reference trajectories determined by each methodology.